Celia bought 12 apples, ate two of them, and sold the rest at .20 more per apple than she paid. Her total profit was $1.00. How much did she sell each apple for?
Answer by stanbon(75887)

Celia bought 12 apples, ate two of them, and sold the rest at .20 more per apple than she paid. Her total profit was $1.00. How much did she sell each apple for?
------------------------
Let her buying price be "x" cents per apple.
Then her selling price is "x+20" cents per apple.
------------------------
Equation:
12 apple Buying price = 12x cents
10 apple Selling price = 10(x+20)
-------------------------
Equation:
selling price - buying price = 100 cents
10(x+20) - 12x = 100
10x + 200 -12x = 100
-2x = -100
x = 50 cents (buying price per apple)
x+20 = 70 cents (selling price per apple)
===============================================
Cheers,
Stan H.